OpenStudy (anonymous):

When a wire is made smaller, the resistance increases. Which happens to the electric current?

OpenStudy (akashdeepdeb):

Do you have options? Ok, even if you don't. \[R = \frac{\rho l}{A}\] Now if \(l\) decreases, definitely \(R\) will also decrease. [Because they are proportional] We know that, \(V = IR\) [From ohm's law] Thus, \(R = \frac{V}{I}\) And therefore, \[\frac{V}{I} = \frac{\rho l}{A}\] When \(V, \rho,~~\& ~~A\) are kept constant. I is inversely proportional to length and hence, when length decreases, electric current increases! PS I think you mean to say when wire length decreases, resistance decreases!